News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 4 May 2001 10.30am EDT

Last week the observing schedule was halted due to high radiation associated with solar flare activity. The loads were halted on Apr 28 at 12:52am EDT through a ground command to execute the SI Safing SCS 107. The projected orbital radiation dose was significantly higher than the threshold allowed for ACIS. The loads were rebuilt and resumed on Apr 29 at 12:40am EDT once the radiation levels had subsided to an acceptable level. Observations of Arp 270 and 3C 228 were impacted and will be rescheduled in a future load.

A key real-time activity was completed on May 2 with the successful uplink and on-orbit test of the Objective Transmission Grating (OTG) flight software patch. The patch allows once more for autonomous operation of the Low Energy Transmission Grating from the daily load. LETG moves have been made with real-time hardware commanding since Aug 2000 to work around anomalous HETG retract limit switch indications that inhibit software commanded LETG motion. In the case of HETG retractions, the new software patch uses data from the potentiometer and grating motion timing to determine the retracted state. Operation of the LETG from the daily loads will commence following approval of the associated ground operations products.

The planned observations for the coming week are shown below and is dominated by HETG observations.

Radiation Belts                      May 6  
Mkn 766                ACIS-S/HETG
Hydra A                ACIS-S/HETG   May 8  
NGC 1399               ACIS-S/HETG
SS 433                 ACIS-S/HETG
WR140                  ACIS-S HETG
Radiation Belts                      May 9  
Centaurus A            ACIS-S/HETG
SS 433                 ACIS-S/HETG   May 10 
RXJ 1347.5-1145        ACIS-S/HETG
Radiation Belts                      May 11 
SS 433                 ACIS-S/HETG   May 12 
3C396 (G39.2-0.3)      ACIS-S
B2 2236+35             ACIS-S        May 13


All spacecraft subsystems continued to operate nominally.

Return to Status Reports